OpenStudy (anonymous):

Write in the form a(x-h)^2...

OpenStudy (anonymous):

Quadratic Equation, write in the form of a(x-h)^2. \[6x^2+36x+16\] Find a= h= k=

OpenStudy (anonymous):

a(x-h)^2 =a(x^2-2xh+h^2) 6x^2+36x+16 =6(x^2+6x+16/6) now we are asked to find a=? h=? and k=? but k isn't mentioned anwhere in the question - so not clear what we need to do with it .